A 0.250g sample of unknown acid requires 89.3 mL of 0.140 mol/L NaOH for complete neutralization. What is the molar mass of acid?


1
Expert's answer
2021-10-10T10:16:48-0400

number of moles of NaOH = "0.140\textsf{M} ×\dfrac{89.3}{10^3}\textsf{ L}" = "0.0125\textsf{ mol}"


Assuming X is monovalent

"HX+ NaOH \to NaX + H_2O"


This implies that one mole of the acid combines with one mole of the base,

"\therefore" 0.0125 moles of the acid combines with 0.0125 moles of the base


0.0125 moles = 0.250g

1 mole = xg

x = "\dfrac{0.250}{0.0125}" = "20\textsf{ g/mol}"